# RiboFlow: Conditional De Novo RNA Sequence-Structure Co-Design via Synergistic Flow Matching

### v2

@@ 5 Experiments
Table 1: Evaluation for Unconditional RNA Generation.
Method %Validity($\uparrow$) Diversity($\uparrow$) Novelty($\downarrow$)
@@
RiboFlow-Codesign
step_50 37.0 0.550 0.540
step_100 34.3 0.545 0.577
Figure 3: The probability distribution statistics of generated data and real data in bond distances and bond angles.
Pretraining Dataset. To establish foundational priors for RNA structural validity, we pre-train RiboFlow on RNAsolo [^1] https://rnasolo.cs.put.poznan.pl (Adamczyk et al., 2022), a curated database of single-stranded RNA 3D structures. We filter entries to those with resolution $\leq$ 4 Å (as of December 2024) and retain sequences between 30–200 nucleotides to balance structural diversity with computational feasibility, yielding 7,154 high-quality training samples. This length range reflects typical functional RNA motifs while accommodating GPU memory constraints. Single-stranded RNAs are prioritized to focus on unpaired regions critical for ligand binding.
RNA-ligand Interaction Dataset. We introduce RiboBind, the largest standardized dataset for RNA-small molecule interactions, addressing data scarcity in ligand-conditioned design. We discuss dataset construction and perform statistical analysis in the Appendix A.1. For sequences of more than 200 nucleotides, we perform dynamic cropping and truncate regions distal to the ligand-binding pocket, preserving interaction sites while maximizing data utility (details discussed in Appendix A.2). The dataset after data augmentation contains 4,445 RNA-ligand pairs.
Dataset Splits. To evaluate generalization, we partition RiboBind into: (1) Standard Evaluation Set: Contains 66 RNA-ligand pairs derived from the 20 most frequent ligands. RNA sequences are clustered at 50% sequence similarity using MMseqs2 (Steinegger & Söding, 2017), with cluster centroids selected to ensure structural diversity. (2) Few-shot Set: Includes 15 pairs with ligands appearing only once in training, testing low-data generalization. Both sets exclude training ligands to prevent leakage.
Evaluation Metrics. We evaluate our experiment using metrics for both generation quality and binding capability. For generation quality, we focus on four metrics: Validity (val.), assessed by inverse-folding each generated backbone with gRNAde (Joshi et al., 2024) and predicting its structure using RhoFold (Shen et al., 2024) with 8 generated sequences. The validity is determined by a self-consistency TM-score (scTM) between the predicted and original backbone at the $C4^{\prime}$ level, with $\text{scTM}\geq 0.45$ indicating a valid backbone; Diversity (div.), measured by the proportion of unique structural clusters (identified by qTMclust (Zhang et al., 2022)) among valid samples, Structures with a TM-score $\geq 0.45$ are considered similar, reflecting the structural variability of generated samples; and Novelty (nov.), evaluated using US-align (Zhang et al., 2022) to compare the structure of valid backbones to the training distribution, with higher novelty indicating greater structural divergence. Sequence Recovery (SR.), quantified as the percentage of correctly recovered nucleotides in the co-designed sequence relative to the ground-truth RNA.
For binding capability, we assess the complex binding capability by four metrics: VinaScore (vina.), the binding free energy (in kcal/mol) between RNA and ligand predicted by GNINA (Dunn et al., 2024), with lower values signifying stronger binding; PoseScore (pose.), a measure of the ligand’s binding pose quality predicted by GNINA, and is defined as the proportion of binding pose probability exceeds the threshold of 0.6; Affinity, predicted binding affinity by GNINA, with higher values indicating stronger binding; and AFScore (AF.), Alphafold3’s overall prediction confidence for the complex, incorporating structural confidences (pTM and ipTM), clash penalties, and considerations for disordered regions, where higher scores are better.
Figure 4: (Left): Performance of two pretraining tasks across different pre-training steps. (Right): Sampling time between RiboFlow (with ligand ARG) and RNA-FrameFlow (without ligand) for RNAs of varying lengths.
Baselines. To the best of our knowledge, there is currently no RNA design model specifically targeting small molecules. RNAFlow (Nori & Jin, 2024) relies on RoseTTAFoldNA (Baek et al., 2024) to design RNA structures from predicted protein-RNA complex structures, which cannot predict RNA-ligand complex conformations. MMDiff (Morehead et al., 2023) is also limited to protein-RNA generation. Meanwhile, LigandMPNN (Dauparas et al., 2023) can only generate proteins for specific ligands. To this end, we choose RNA-FrameFlow (Anand et al., 2024) for comparison as it is a general RNA structure generation model, despite not being constrained by ligands. For further comparison, we also include a baseline where sequences are randomly generated and folded using RhoFold.
Table 2: Comparison of RNA generation under different conditions.
Vina. ()($\downarrow$)( ↓ ) Affinity ()($\uparrow$)( ↑ ) %Pose. ()($\uparrow$)( ↑ ) %AF. ()($\uparrow$)( ↑ ) %Val.()($\uparrow$)( ↑ ) Div.()($\uparrow$)( ↑ ) Nov.()($\downarrow$)( ↓ )
Top10 Median Top10 Median
@@
Random -2.76 -2.55 2.34 2.31 4.20 3.11 - - -
RNA-FrameFlow -4.15 -4.01 3.12 3.02 59.6 17.4 22.5 0.288 0.584
RiboFlow -4.25 -4.12 3.54 3.28 69.4 22.8 10.8 0.226 0.640
+ Pre -4.39 -4.31 3.56 3.34 71.8 48.4 30.8 0.354 0.514
+ Crop -4.38 -4.27 3.59 3.30 70.3 34.7 12.1 0.293 0.603
+ Pre + Crop -4.30 -4.18 3.61 3.35 78.5 50.1 24.2 0.376 0.534
@@
Random -2.91 -2.82 2.41 2.24 5.24 5.25 - - -
RNA-FrameFlow -4.27 -4.15 3.32 3.14 55.7 22.7 22.7 0.545 0.574
RiboFlow -4.61 -4.48 3.79 3.55 65.5 25.7 10.8 0.517 0.669
+ Pre -4.67 -4.55 3.92 3.70 73.6 51.2 35.8 0.575 0.519
+ Crop -4.63 -4.56 3.94 3.66 66.9 38.9 12.7 0.522 0.613
+ Pre + Crop -4.58 -4.43 3.82 3.60 74.6 54.8 25.6 0.589 0.527
@@
Random -3.32 -3.13 2.53 2.51 6.00 3.54 - - -
RNA-FrameFlow -4.51 -4.48 3.61 3.54 65.3 22.3 23.9 0.556 0.581
RiboFlow -5.01 -4.82 4.21 3.97 62.3 15.4 11.2 0.522 0.656
+ Pre -5.12 -5.07 4.40 4.11 72.6 45.7 34.1 0.561 0.537
+ Crop -5.10 -4.99 4.29 4.02 66.7 27.9 12.3 0.546 0.609
+ Pre + Crop -4.96 -4.84 4.31 4.03 73.8 39.8 26.1 0.559 0.540
RiboFlow-T -5.07 -4.94 4.42 4.05 75.1 46.2 22.4 0.550 0.524
5.1 Unconditional RNA Generation
Setup. We evaluate the generative capabilities of models without small molecule information conditioned. Meanwhile, to validate the effectiveness of the co-design pre-training strategy proposed in the paper, we design two different pre-training tasks: (i) utilizing only the RNA structural backbone frames and torsion angle loss (RiboFlow-Structure), and (ii) incorporating nucleotide type loss simultaneously (RiboFlow-Codesign). For each model, we set the RNA generation length within the range of [50, 150], sampling at intervals of 20, and generate 50 samples for each length to thoroughly cover the model’s sampling space.
Results. The experimental results are shown in Table 1, with local structural metrics of RNA generated by the RiboFlow-Codesign depicted in Figure 3. Additionally, we provide a detailed comparison of model performance under different pre-training steps in Table S1. We can draw the following conclusions: (i) The validity of RNAs generated by RiboFlow-Codesign has shown significant improvement compared to the other baselines. (ii) Due to the addition of sequence feature constraints, the diversity of RNAs generated by the Codesign model has slightly decreased compared to RiboFlow-Structure. (iii) Increasing the sampling steps can enhance the validity of RNA structures to a certain extent but will reduce the diversity of the model. Besides, it can also be noted that the structures generated by RiboFlow-Codesign are close to the real structures in terms of bond lengths and torsion angles in Table 1, which further proves the effectiveness of the co-design pre-training strategy.
@@ 5.2 RNA Generation Guided by Ligand Structure
We evaluate the performance of models conditioned on ligand-bound structures and further examine the impact of incorporating the ground-truth RNA sequence length as a prompt. Hence, we design two protocols: (a) using the true RNA sequence length for generation, and (b) employing uniform sampling of RNA lengths within a predefined range. Both protocols are tested on the standard evaluation set. Additionally, we expand our analysis to include the few-shot evaluation set (detailed in the Appendix Table S2) to assess the model’s performance in low-resource scenarios.
Table 3: Sequence recovery and structural RMSD comparisons with ground-truth RNA using ligand-bound structures and true sequence lengths.
@@
RNA-FrameFlow 32.2 8.92
RiboFlow 37.9 9.66
+Pre 47.4 5.94
+Crop 40.6 7.03
+Pre+Crop 43.8 7.34
Setup. We include three types of model variants in addition to the original model: (i) RiboFlow: The baseline model trained on the RiboBind; (ii) +CROP: Model trained on the RiboBind augmented with dynamic cropping strategy; (iii) +PRE: Model trained on the RiboBind using codesign pretrained weights; (iv) +PRE+CROP: Model trained on RiboBind augmented with dynamic cropping strategy using codesign pretrained weights. The ligand-bound structures are provided to models as guidance. For protocol (a), 50 RNAs of actual length are sampled for each RNA-ligand pair. As for protocol (b), the range is defined as [50,150], sampling at intervals of 20, and generated 50 samples for each length, resulting in a total of 300 RNAs. These candidate RNAs are subsequently evaluated based on their structural validity and binding capability with the ligand.
Results. We first analyze the experimental results provided with the ground-truth RNA lengths, as shown in Table 2 in the gray-shaded section and Table 3, with best results in bold and the second-best results underlined. Furthermore, we provide detailed experimental results for 24 randomly selected RNA-ligand pairs presented in Figures S7, showcasing the vina and posescore with actual values as reference. Several interesting conclusions can be summarized: (i) The diversity for all models are significantly lower than those observed in Table 1, which indicates that constraining the RNA length compresses the sequence sampling space, leading to a decrease in diversity; (ii) Models incorporating co-design pre-training strategy (PRE and PRE-CROP) demonstrate a substantial advantage in structural validity and binding capability. Besides, we can also observe that these two models also perform excellently in terms of sequence recovery rate and RMSD metrics in Table 3; (iii) Compared to the codesign-pretraining, there is no significant improvement in structural validity under the dynamic cropping strategy. This may be because the trimmed RNA fragments may not fold into the real RNA structure, thus the model cannot effectively learn the true structural distribution. However, data augmentation expands the training data, allowing the model to learn a more diverse set of RNA-ligand docking patterns and improve the designed RNA binding capability.
Subsequently, we analyze the experimental results when length is sampling from the pre-defined ranges in Table 2 highlighted in the blue-shaded section. Most experimental conclusions are similar to the previous ones, but due to the increase in the length sampling space, models can explore more potential binding patterns, and the results of all model variants on Vina, Affinity, and AFScore increase largely.
We conduct similar experiments on few-shot set, and experimental results can be referred to the Appendix B.3.
@@ 5.3 RNA Generation Guided by Ligand SMILES
Setup. We employ RDKit to generate 3D ligand conformers from SMILES, with subsequent optimization using the MMFF force field (Tosco et al., 2014). RNA lengths are sampled at intervals of 20 within the range [50, 150], with 50 candidate structures generated per length. Moreover, we introduce RiboFlow-T, a variant that translates the coordinate system to the ligand centroid during the training of RiboFlow-PRE-CROP. This modification aims to minimize the impact of ligand spatial positioning on RNA-ligand complex modeling. The nucleotide structure frames are subsequently constructed relative to this transformed coordinates.
Results. The experimental results are shown in Table 2 highlighted in the pink-shaded section, which reveals the following insights: RiboFlow-T achieves the highest docking pose score, indicating that re-centering on the ligand’s centroid during training improves docking pattern reliability. However, this approach introduces inconsistencies with the RNA-centered coordinate system used during pretraining, reducing RNA structure validity.
We also leverage RiboFlow to de novo design RNA conditioned on ligand A2F and GNG, with the experimental results presented in the Figure 5 and Appendix B.4.
Figure 5: De novo design of an RNA corresponding to ligand A2F. Left: The ground-truth RNA-ligand complex (PDB ID: 3GOT). Right: RNA backbone designed by RiboFlow, with ligand docking by GNINA. The designed structure achieves a TM-score of 0.47 and a Sequence Recovery (SR) of 41.8% compared to the ground-truth complex.

### v3

@@ 5 Experiments
Pretraining Dataset. To establish foundational priors for RNA structural validity, we pre-train RiboFlow on RNAsolo adamczyk2022rnasolo , a curated database of single-stranded RNA 3D structures. We filter entries to those with resolution $\leq$ 4 Å (as of December 2024) and retain sequences between 30–200 nucleotides to balance structural diversity with computational feasibility, yielding 7,154 high-quality training samples. This length range reflects typical functional RNA motifs while accommodating GPU memory constraints.
RNA-ligand Interaction Dataset. We introduce RiboBind, a large standardized dataset for RNA-small molecule interactions, addressing data scarcity in ligand-conditioned design. We discuss dataset construction and perform statistical analysis in the Appendix 7.1. We perform dynamic cropping and truncate regions distal to the ligand-binding pocket, preserving interaction sites while maximizing data utility (details discussed in Appendix 7.3). Meanwhile, we have conducted a comprehensive and objective comparison of existing RNA-ligand datasets, which can be referred to in the Appendix 7.2.
Dataset Splits. To evaluate model generalization, we employ the following partitioning strategies for RiboBind: (1) Sequence-based Evaluation: RNA sequences are clustered at a 50% identity threshold using MMseqs2 steinegger2017mmseqs2 . A test set is formed from the cluster centroids, comprising 66 RNA-ligand pairs which include 20 of the most frequently observed ligands. All non-centroid sequences constitute the corresponding training set. (2) Structure-based Evaluation: Following the methodology of gRNAde joshi2024grnade , RNA structures are clustered using US-align zhang2022us with a TM-score threshold of 0.45, yielding 277 distinct structural classes. These classes are then partitioned into a training set (249 classes) and a test set (28 classes) at a 9:1 ratio. The final test dataset for this evaluation consists of 28 RNA-ligand pairs, each formed by one representative structure from a held-out test class and its corresponding ligand. (3) Few-shot Evaluation: Contains 15 RNA-ligand pairs involving ligands that appear only once in the RiboBind dataset, designed to evaluate low-resource generalization. In the main text, we primarily introduce sequence-based evaluation, while the appendix provides a detailed analysis for structure-based evaluation and few-shot evaluation.
Evaluation Metrics. We evaluate our experiment using metrics for both generation quality and binding capability. For generation quality, we focus on four metrics: Validity (val.), assessed by inverse-folding each generated backbone with gRNAde joshi2024grnade and predicting its structure using RhoFold shen2024accurate with 8 generated sequences. The validity is determined by a self-consistency TM-score (scTM) between the predicted and original backbone at the $C4^{\prime}$ level, with $\text{scTM}\geq 0.45$ indicating a valid backbone; Diversity (div.), measured by the proportion of unique structural clusters (identified by qTMclust zhang2022us ) among valid samples, Structures with a TM-score $\geq 0.45$ are considered similar, reflecting the structural variability of generated samples; and Novelty (nov.), evaluated using US-align zhang2022us to compare the structure of valid backbones to the training distribution, with higher novelty indicating greater structural divergence. Sequence Recovery (SR.), quantified as the percentage of correctly recovered nucleotides in the co-designed sequence relative to the ground-truth RNA.
For binding capability, we assess the complex binding capability in-silico by three metrics: AFScore (AF.), Alphafold3’s overall prediction confidence for the complex, incorporating structural confidences (pTM and ipTM), clash penalties, and considerations for disordered regions, where higher scores are better; GerNAScore (measured by GerNA-Bind  [^1] https://github.com/GENTEL-lab/GerNA-Bind), A recently proposed model for RNA–ligand affinity prediction xia2025gerna , where higher scores correspond to stronger predicted binding affinity. VinaScore (vina.), the binding free energy (in kcal/mol) between RNA and ligand predicted by AutoDock Vina trott2010autodock , with lower values signifying stronger binding capacity.
Baselines. To the best of our knowledge, there is currently no RNA design model specifically targeting small molecules. RNAFlow nori2024rnaflow relies on RoseTTAFoldNA baek2024accurate to design RNA structures from predicted protein-RNA complex structures, which cannot predict RNA-ligand complex conformations. MMDiff morehead2023towards is also limited to protein-RNA generation. LigandMPNN dauparas2023atomic can only generate proteins for specific ligands. To this end, we choose RNA-FrameFlow anand2024rnaframeflow for comparison as it is a general RNA structure generation model, despite not being constrained by ligands. For further comparison, we also include a baseline where sequences are randomly generated and folded using RhoFold.
5.1 Unconditional RNA Generation
Setup. We first evaluate the generative capabilities of each model without conditioning on ligand information. In addition to using the original RNAFrame-Flow model, we assess the effectiveness of the proposed co-design pre-training strategy by comparing: (i) RNA-FrameFlow-R, a retrained version of RNA-FrameFlow using the new pre-training dataset with RNA backbone frames and torsion angle loss; and(ii) RiboFlow, the proposed model trained with the sequence-structure co-design loss. For each model, we set the generation length in the range of [50, 150], sampling at intervals of 20, and generate 100 samples for each length to cover the model’s sampling space.
Table 1: Unconditional RNA generation.
Method %Val.($\uparrow$) Div.($\uparrow$) Nov.($\downarrow$)
@@ 5.1 Unconditional RNA Generation
RiboFlow
step_50 34.7 0.550 0.540
step_100 31.9 0.545 0.577
Results. The experimental results are presented in Table 1, where step_50 denotes a sampling step of 50. A more detailed comparison of model performance under varying pre-training steps is provided in Table 4. The local structural properties of RNAs generated by RiboFlow and RNAFrame-Flow-R are illustrated in Figure 8. We can draw the following conclusions: (i) The validity of RNAs generated by RiboFlow has shown significant improvement compared to the other baselines. (ii) Due to the addition of sequence feature constraints, the diversity of RNAs generated by the RiboFlow has slightly decreased compared to RNA-FrameFlow-R. (iii) Increasing the sampling steps can enhance the validity of RNA structures to a certain extent. Besides, it can also be noted that the structures generated by RiboFlow are close to the real structures in terms of bond lengths and torsion angles in Figure 8.
Table 2: Comparison of RNA generation with sequence-based evaluation.
Vina. $(\downarrow)$ GerNA. $(\uparrow)$ %AF. $(\uparrow)$ %Val.()($\uparrow$) Div.()($\uparrow$) Nov.()($\downarrow$)
Top10 Median
@@
Random -2.76 -2.55 0.101 3.11 - - -
RNA-FrameFlow -4.15 -4.01 0.203 17.4 22.5 0.288 0.584
RiboFlow -4.25 -4.12 0.283 22.8 10.8 0.226 0.640
+ Pre -4.39 -4.31 0.437 48.4 26.8 0.354 0.514
+ Crop -4.38 -4.27 0.402 34.7 12.1 0.293 0.603
+ Pre + Crop -4.30 -4.18 0.411 50.1 24.2 0.376 0.534
@@
Random -2.91 -2.82 0.113 5.25 - - -
RNA-FrameFlow -4.27 -4.15 0.209 22.7 22.7 0.545 0.574
RiboFlow -4.61 -4.48 0.297 25.7 10.8 0.517 0.669
+ Pre -4.67 -4.55 0.467 51.2 25.9 0.575 0.519
+ Crop -4.63 -4.56 0.456 38.9 12.7 0.522 0.613
+ Pre + Crop -4.58 -4.43 0.448 54.8 25.6 0.589 0.527
@@
Random -3.32 -3.13 0.107 3.54 - - -
RNA-FrameFlow -4.51 -4.48 0.200 22.3 23.9 0.556 0.581
RiboFlow -5.01 -4.82 0.304 15.4 11.2 0.522 0.656
+ Pre -5.12 -5.07 0.469 45.7 28.1 0.561 0.537
+ Crop -5.10 -4.99 0.451 27.9 12.3 0.546 0.609
+ Pre + Crop -4.96 -4.84 0.460 39.8 26.1 0.559 0.540
RiboFlow-T -5.07 -4.94 0.472 46.2 22.4 0.550 0.524
@@ 5.2 RNA Generation Guided by Ligand Structure
We evaluate the performance of models conditioned on ligand-bound structures and further examine the impact of incorporating the ground-truth RNA sequence length as a prompt. Hence, we design two protocols: (a) using the true RNA sequence length for generation, and (b) employing uniform sampling of RNA lengths within a predefined range. Both protocols are tested on the sequence-based evaluation set. Additionally, we expand our analysis to include the structure-based and few-shot evaluation set (detailed in the Appendix Table 6 and Table 7) to assess the model’s performance.
Table 3: Sequence recovery and structural RMSD comparisons with ground-truth RNA using ligand-bound structures and true sequence lengths on the test set.
@@
RNA-FrameFlow 28.2 12.32
RiboFlow 30.9 11.66
+Pre 33.4 9.43
+Crop 31.2 10.32
+Pre+Crop 31.8 11.64
Setup. We include three types of model variants in addition to the original model: (i) RiboFlow: The baseline model trained on the RiboBind; (ii) +CROP: Model trained on the RiboBind augmented with dynamic cropping strategy; (iii) +PRE: Model trained on the RiboBind using codesign pretrained weights; (iv) +PRE+CROP: Model trained on RiboBind augmented with dynamic cropping strategy using codesign pretrained weights. The ligand-bound structures are provided to models as guidance. For protocol (a), 100 RNAs of actual length are sampled for each RNA-ligand pair. As for protocol (b), the range is defined as [50,150], sampling at intervals of 20, and generated 100 samples for each length. These candidate RNAs are subsequently evaluated based on their structural validity and binding capability with the ligand.
Results. We first analyze the experimental results provided with the ground-truth RNA lengths, as shown in Table 2 in the gray-shaded section and Table 3, with best results in bold and the second-best results underlined. Furthermore, we provide detailed experimental results for 24 randomly selected RNA-ligand pairs presented in Figures 12, showcasing the vina with actual values as reference. Several interesting conclusions can be summarized: (i) The diversity for all models are significantly lower than those observed in Table 1, which indicates that constraining the RNA length compresses the sequence sampling space, leading to a decrease in diversity; (ii) Models incorporating co-design pre-training strategy (PRE and PRE-CROP) demonstrate a substantial advantage in structural validity and binding capability. Besides, we can also observe that these two models also perform excellently in terms of sequence recovery rate and RMSD metrics in Table 3; (iii) Compared to the codesign-pretraining, there is no significant improvement in structural validity under the dynamic cropping strategy. This may be because the trimmed RNA fragments may not fold into the real RNA structure, thus the model cannot effectively learn the true structural distribution. However, data augmentation expands the training data, allowing the model to learn a more diverse set of RNA-ligand docking patterns and improve the designed RNA binding capability.
Subsequently, we analyze the experimental results when length is sampling from the pre-defined ranges in Table 2 highlighted in the blue-shaded section. Most experimental conclusions are similar to the previous ones, but due to the increase in the length sampling space, models can explore more potential binding patterns, and the results of Vina, GerNA and AFScore increase largely.
We conduct similar experiments on structure-based evaluation and few-shot evaluation, and experimental results can be referred to the Appendix 8.5 and Appendix 8.6.
@@ 5.3 RNA Generation Guided by Ligand SMILES
Setup. We use RDKit to generate 3D ligand conformers from SMILES, with optimization using the MMFF force field tosco2014bringing . RNA lengths are sampled at intervals of 20 in the range [50, 150], with 200 structures generated per length. Moreover, we introduce RiboFlow-T, a variant that translates the coordinate system to the ligand centroid during the training of RiboFlow-PRE-CROP. This modification aims to minimize the impact of ligand spatial positioning on RNA-ligand complex modeling. The nucleotide structure frames are subsequently constructed relative to this transformed coordinates.
Figure 3: De novo design of an RNA corresponding to ligand A2F. Left: The ground-truth RNA-ligand complex (PDB ID: 3GOT). Right: RNA backbone designed by RiboFlow. The designed structure achieves a TM-score of 0.47 and a Sequence Recovery (SR) of 41.8% compared to the ground-truth complex.
Results. The experimental results are shown in Table 2 highlighted in the pink-shaded section, which reveals the following insights: RiboFlow-T achieves the highest AFScore and GerNA score, indicating that re-centering on the ligand’s centroid during training improves docking pattern reliability. However, this approach introduces inconsistencies with the RNA-centered coordinate system used during pretraining, reducing RNA structure validity. Meanwhile, we observe that when only the SMILES is provided, the predicted RNA-binding affinity is generally higher than that obtained using the true ligand structure. We speculate that this is because the RNA molecule is designed based on the ligand structure generated from SMILES, resulting in a standardized and reproducible input that enhances model robustness. However, the true ligand binding conformation may deviate from the lowest-energy state due to conformational constraints within the complex. In addition, structural variations arising from different sources can introduce input heterogeneity, potentially affecting the model’s generalization ability.
We also leverage RiboFlow to de novo design RNA conditioned on ligand A2F and GNG, with the experimental results presented in the Figure 3 and Appendix $8.$